OpenStudy (anonymous):

if [ -4 5 -2 5/2] find a non-zero 2 by 2 zero matrix X, such that A.X=O_2x2

OpenStudy (amistre64):

is O spose to be anything special? other than a 2x2 matrix?

OpenStudy (amistre64):

and I take it the given matrix is A:

OpenStudy (anonymous):

\[\left[\begin{matrix}-4 & 5 \\ -2 & 5/2\end{matrix}\right]\]

OpenStudy (amistre64):

find a nonzero, zero matrix?

OpenStudy (amistre64):

A.X = 0 is what i think im reading, such that X is not a 0

OpenStudy (amistre64):

so they need to be orthogonal at best

OpenStudy (amistre64):

\[z_{11}=-4x_{11}+5x_{21}=0\] \[z_{21}=-2x_{11}+\frac{5}{2}x_{21}=0\] \[z_{12}=-4x_{12}+5x_{22}=0\] \[z_{22}=-2x_{12}+\frac{5}{2}_{22}=0\] would be our corresponding equations maybe

OpenStudy (anonymous):

A.X= to the letter O and not 0

OpenStudy (amistre64):

the name is unimportant methinks, but its spose to zero out

OpenStudy (amistre64):

-4x + 5y = 0 4x - 5y = 0 ----------- i dont thinks its doable but i could be wrong

OpenStudy (amistre64):

detA = 0 so its dependant for starters, i think that throws a monkeywrench in the plans

OpenStudy (amistre64):

or maybe, since its just one vector, we can flip and negate the entries to get a perp vector; hmmm \[\begin{pmatrix}-4&5\\-2&\frac{5}{2}\end{pmatrix}.\begin{pmatrix}2&-\frac{5}{2}\\-4&5\end{pmatrix}\] maybe?

OpenStudy (amistre64):

that doesnt work either

OpenStudy (amistre64):

i was on the right track :)